OpenStudy (anonymous):

Which is a factor of 8x2 – 12x – 8 ? x + 1 x – 4 2x – 2 2x + 1

OpenStudy (anonymous):

(x-2) (2x+1)

OpenStudy (anonymous):

\(8x^2–12x –8 =4(x-2)(2x+1)\)

OpenStudy (nowhereman):

You can factorize a polynomial by finding its roots. If it has the roots x_i you can write is as \[\text{constant}\cdot (x-x_0)\cdot(x-x_1)\cdot\cdots\cdot(x-x_n)\]

OpenStudy (anonymous):

So the answer from the problem statement is: 2x + 1